OpenStudy (thomas5267):

A homework question for a student (not me) preparing for Chinese Gao Kao. Let \(f(x)\) be a function with domain \(\mathbb{R}\). For all \(x_1,x_2\in \mathbb{R}\), there exist a real L such that \(|f(x_1)-f(x_2)|\leq L|x_1-x_2|\). For \(0

OpenStudy (thomas5267):

If I am not mistaken, f(x) is Lipschitz continuous. According to wikipedia, f(x) is also a contraction.

OpenStudy (thomas5267):

The first part is simple. The second part is extraordinarily hard. And this is a homework question for students preparing for Chinese Gao Kao (A Level equivalent)...

OpenStudy (kainui):

Just curious, what is Lipschitz continuous?

OpenStudy (thomas5267):

A function \(f(x)\) with \(|f(x_1)-f(x_2)|\leq L|x_1-x_2|\) for all \(x_1,x_2\) in the domain of f. L is a fixed positive real number.

OpenStudy (thomas5267):

The second part reminds me of Cesaro summation.

OpenStudy (kainui):

I totally see that Cesaro sum thing. Here's an idea to try to alter the sum a little. \[A_{n+1} = \frac{1}{n+1} \sum_{k=1}^{n+1} a_k = \frac{1}{n+1}\left(a_{n+1}+n* \frac{1}{n} \sum_{k=1}^{n} a_k \right)\] \[A_{n+1} = \frac{1}{n+1}\left(a_{n+1}+nA_n \right) = \frac{a_{n+1}}{n+1} + \frac{n}{n+1} A_n\] \[|A_{n+1}-A_n| = \left|\frac{a_{n+1}}{n+1} + \left(\frac{n}{n+1} -1 \right)A_n \right| = \frac{|a_{n+1}-A_n|}{n+1} \]

OpenStudy (thomas5267):

If I can prove \(a_{k+1}\leq a_k\) then the second part follows.

OpenStudy (thomas5267):

@Kainui Did you not try the first part or couldn't be bothered to type out the answer?

OpenStudy (kainui):

I didn't look at it cause you said it was simple so I assumed you solved it already.

OpenStudy (thomas5267):

Yes I solved it already because it is really simple.

OpenStudy (kainui):

I don't know how to solve any of this though myself, so I have no clue. I don't really care for analysis personally lol

OpenStudy (thomas5267):

For part 1, look at \(|a_k-a_{k+1}|\), then look at the definition of \(a_{k+1}\).

OpenStudy (thomas5267):

Victory is close... Part 1 \[ \begin{align*} &\phantom{{}={}}|a_k-a_{k+1}|\\ &=|f(a_{k-1})-f(a_k)|&\text{Definition of }a_k\\ &\leq L|a_{k-1}-a_k|&\text{Property of }f(x)\\ &\leq L^2|a_{k-2}-a_{k-1}|&\text{Do it again}\\ &\leq L^{k-1}|a_1-a_2| \end{align*} \] \[ \begin{align*} &\phantom{{}={}}\sum_{k=1}^n|a_k-a_{k+1}|\\ &\leq \sum_{k=1}^n L^{k-1}|a_1-a_2|\\ &=|a_1-a_2|\sum_{k=0}^{n-1} L^{k} &\text{Index shift}\\ &<|a_1-a_2|\sum_{k=0}^\infty L^{k}\\ &=\frac{1}{1-L}|a_1-a_2|&0<L<1 \end{align*} \] Part 2 \[ \begin{align*} |a_k-a_{k+1}|&\leq L^{k-1}|a_1-a_2|\\ |a_{k-1}-a_k|&\leq L^{k-2}|a_1-a_2|\\ |a_k-a_{k+1}|+|a_{k-1}-a_k|&\leq |a_1-a_2|(L^{k-1}+L^{k-2})\\ |a_{k-1}-a_{k+1}|&\leq |a_1-a_2|(L^{k-2}+L^{k-3})\text{ Triangle inequality}\\ |a_i-a_j|&\leq|a_1-a_2|\left(\sum_{s=i-1}^{j-2} L^s\right)\text{ in general}\\ |a_i-a_j|&<|a_1-a_2|\left(\sum_{s=1}^\infty L^s\right)\text{ min. }i=1\\ |a_i-a_j|&<\frac{1}{1-L}|a_1-a_2| \end{align*} \] \[ \begin{align*} &\phantom{{}={}}|A_k-A_{k+1}|\\ &=\left|\frac{1}{k}\sum_{j=1}^k a_j-\frac{1}{k+1}\sum_{j=1}^{k+1} a_j \right|\\ &=\frac{1}{k(k+1)}\left|(k+1)\sum_{j=1}^k a_j-k\sum_{j=1}^k+1 a_j \right|\\ &=\frac{1}{k(k+1)}\left|\sum_{j=1}^k (a_j)-ka_{k+1}\right|\\ &=\frac{1}{k(k+1)}\left|\sum_{j=1}^k (a_j-a_{k+1})\right|\\ &\leq\frac{1}{k(k+1)}\sum_{j=1}^k \left|a_j-a_{k+1}\right|&\text{Triangle inequality}\\ &\leq\frac{1}{k(k+1)}\sum_{j=1}^k \left(\frac{1}{1-L}|a_1-a_2|\right)\\ &=\frac{1}{k(k+1)}\left(\frac{k}{1-L}|a_1-a_2|\right)\\ &=\frac{1}{(k+1)}\left(\frac{1}{1-L}\right)|a_1-a_2| \end{align*} \]

OpenStudy (thomas5267):

@ganeshie8 Help! Finished part 1 but part 2 just so hard!

Parth (parthkohli):

what

OpenStudy (thomas5267):

Is it too confusing or what?

OpenStudy (thomas5267):

The question is in Chinese so I have to translate it into English.